# Overcoming Reward Model Noise in Instruction-Guided Reinforcement Learning

### v1

@@ A.2 Proof of the Reduction of Convergence Rate
    We use $\Pi_{G}$ to represent the set of acceptable policies and use $\Pi_{B}$ as an arbitrary set of “not acceptable” policies that are learned to follow instruction but fail to reach subgoals, i.e., $\Pi_{B}\cap\Pi_{G}=\varnothing$.
@@
    We assume that $|\Pi_{G}|$ and $|\Pi_{B}|$ is a fixed number predefined by the task environment while $\mathbb{E}[G^{\pi}|\pi\in\Pi_{B}]$ is non-zero as false positive rewards are unavoidable in real-world VLMs.
Since the update rule for $Q$-value is a gradient descent on $\|\mathbb{E}[G^{\pi_{\theta}}-Q_{\phi}(s,a)]\|^{2}$ and also we have that $\{\Pi_{B},\Pi_{G},\Pi\setminus(\Pi_{B}\cup\Pi_{G})\}$ is a countable partition of the policy universe $\Pi$, the updated $Q$-value will approach as follows:
$\quad Q_{\phi}(s,a)\rightarrow\mathbb{E}[G^{\pi_{\theta}}]$
$=P(\pi_{\theta}\in\Pi_{B})\cdot\mathbb{E}[G^{\pi_{\theta}}|\pi_{\theta}\in\Pi_{B}]$
$\qquad+P(\pi_{\theta}\in\Pi_{G})\cdot\mathbb{E}[G^{\pi_{\theta}}|\pi_{\theta}\in\Pi_{G}]$
$\qquad+P(\pi_{\theta}\in\Pi\setminus(\Pi_{B}\cup\Pi_{G}))$
$\qquad\cdot\mathbb{E}[G^{\pi_{\theta}}|\pi_{\theta}\in\Pi\setminus(\Pi_{B}\cup\Pi_{G})]$
$=P(\pi_{\theta}\in\Pi_{B})\cdot\mathbb{E}[G^{\pi_{\theta}}|\pi_{\theta}\in\Pi_{B}]$
$\qquad+P(\pi_{\theta}\in\Pi_{G})\cdot\mathbb{E}[G^{\pi_{\theta}}|\pi_{\theta}\in\Pi_{G}]$ (6)
Given that the update rule for policy $\pi$ is the gradient ascent on $Q_{\phi}(s,a)\pi_{\theta}(a|s)$, we have the following:
$\quad\nabla_{\theta}Q_{\phi}(s,a)\pi_{\theta}(a|s)$
$=(\nabla_{\theta}P(\pi_{\theta_{old}}\in\Pi_{B})\cdot\mathbb{E}[G^{\pi_{\theta_{old}}}|\pi_{\theta_{old}}\in\Pi_{B}]\pi_{\theta}(a|s))$
$\qquad+(\nabla_{\theta}P(\pi_{\theta_{old}}\in\Pi_{G})$
$\qquad\cdot\mathbb{E}[G^{\pi_{\theta_{old}}}|\pi_{\theta_{old}}\in\Pi_{G}]\pi_{\theta}(a|s))$
$=P(\Pi_{B})\cdot\mathbb{E}[G^{\Pi_{B}}]\nabla_{\theta}\pi_{\theta}(a|s)$
$\qquad+P(\pi_{G})\cdot\mathbb{E}[G^{\pi_{G}}]\nabla_{\theta}\pi_{\theta}(a|s)$
$=(1-P(\pi_{G})-P(\pi_{\theta}\in\Pi\setminus(\Pi_{B}\cup\Pi_{G})))$
$\qquad\cdot\mathbb{E}[G^{\Pi_{B}}]\nabla_{\theta}\pi_{\theta}(a|s)+P(\pi_{G})\cdot\mathbb{E}[G^{\pi_{G}}]\nabla_{\theta}\pi_{\theta}(a|s)$
$=\textit{const}\cdot\mathbb{E}[G^{\Pi_{B}}]\nabla_{\theta}\pi_{\theta}(a|s)$
$\qquad+(\mathbb{E}[G^{\pi_{G}}]-\mathbb{E}[G^{\Pi_{B}}])P(\pi_{G})\nabla_{\theta}\pi_{\theta}(a|s)$ (7)
Justification of the assumptions.
  • Regarding the non-zero probability of recovering the optimal policy at initialization, it is standard in theoretical analyses to assume a uniform distribution of a random variable at initialization (see Agarwal et al. (2021)). This assumption does not contradict the conclusion about the convergence rate deterioration in Theorem 3.1.
  • In reference to the realizability condition implied by Assumption 2, the expressiveness of the policy class parameterized by $\theta$ is an underlying assumption for deep learning models, supported by the The Universal Approximation Theorem (Hornik et al. 1989).
Observation. Since the goal of the learning agent is to maximize $P(\pi\in\Pi_{G})$ (i.e., to converge to an acceptable policy), we can see that the second term provides the target direction with rate $(\mathbb{E}[G^{\pi\in\Pi_{G}}]-\mathbb{E}[G^{\pi\in\Pi_{B}}])$. Therefore, the ascent rate will decrease when the cumulative rewards from trajectories that cannot reach the goal state (i.e., the false positive rewards) gets higher. Moreover, the first term $\textit{const}\cdot\mathbb{E}[G^{\Pi_{B}}]\nabla_{\theta}\pi_{\theta}(a|s)$ can be regarded as the deviation of target direction. It shows that the level of deviation is also positively proportional to the magnitude of rewards coming from failed trajectories. This theorem follows the intuition that the presence of false positive rewards can slow down the convergence rate of the learning agent.
A.3 Implementation Details of the Experiments
Environment Details
@@ A.3 Implementation Details of the Experiments
We describe each testing environment used in our experiments. More details introduction can be found in on the official project homepage of each benchmark (Chevalier-Boisvert et al. 2018; Bellemare et al. 2013; Hafner 2021).
@@
Instruction-Guide Procedure Details
The VLM-based reward model will have a pointer to the sequence of the instruction sentence, starting at the first sentence. For original models Pixl2R and ELLM-, we follow the setting in there original work where for each instruction sentence, the reward model will have a maximum cap of rewards (2.0) it can assign to the agent in one episode. When the cap is reached, the reward model will move its pointer to the next instruction sentence. For the BiMI reward model, the reward model will move its pointer to the next instruction sentence when the binary signal is triggered.
Finetuning VLM-based Reward Models
In contrast to previous work on Instruction-guided RL where they rely on hand-crafted oracle multimodal reward models, we use actual pretrained VLMs to generate reward signals. 2 VLM backbone models are used in our experiments: 1) CLIP (Radford et al. 2021), pretrained by image-text pairs; and (2) X-CLIP (Ma et al. 2022), pretrained by video-text pairs. In particular, Pixl2R uses CLIP because it only uses the single latest frame as input. In contrast, ELLM- takes a slice of trajectory (i.e., multiple frames) as input, and thus uses either X-CLIP or CLIP with additional RNN encoder as the reward model.
@@
Hyperparameters for Instruction-Guided RL Agents
In the experiments, all methods are implemented based on PPO with same model architecture. The Minigrid and Crafter environments use the same training hyperparameters as the Achievement Distillation paper (Moon et al. 2023). For Montezuma’s Revenge, we found that the performance of the agent was sensitive to the gamma and GAE lambda parameters. To improve the performance of agents in Montezuma’s Revenge, we took two additional steps: (1) normalizing the observation inputs when computing the rewards, and (2) not normalizing the advantage during the GAE calculation. The hyperparameters are shown in the following tables.

### v4

@@ A.2.2 False Positive and the Violation of Pessimistic Property of Heuristics
    for all successor states $s^{\prime}$: $h(s^{\prime})\leq V^{*}(s^{\prime})$ for an arbitrary $s$.
@@
    false positive at current arbitrary state $s$: $V^{*}(s)<h(s)$.
The objective is to show that under the above conditions, we obtain $\max_{a}(\mathcal{B}h)(s,a)<h(s)$.
  1. 1.
    Express the Bellman backup for $h$: $(\mathcal{B}h)(s,a)=R(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)h(s^{\prime})$
  2. 2.
    Express the Bellman version of $V^{*}$: $V^{*}(s)=\max_{a\in\mathcal{A}}\left[R(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)V^{*}(s^{\prime})\right]$.
    Given that $h(s^{\prime})\leq V^{*}(s^{\prime})$ for all $s^{\prime}$, we can expand it to $\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)h(s^{\prime})\leq\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V^{*}(s^{\prime})$. Therefore, we have:
    $R(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)h(s^{\prime})\leq R(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V^{*}(s^{\prime})=Q^{*}(s,a)$ (27)
@@
    Taking the maximum over actions to both sides, we have:
    $\max_{a}\left(R(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)h(s^{\prime})\right)$ $\leq max_{a}Q^{*}(s,a)$ (28)
    $\max_{a}(\mathcal{B}h)(s,a)$ $\leq V^{*}(s)$ (29)
  4. 4.
    Apply false positive condition: Given $V^{*}(s)<h(s)$, substituting into the above inequality, we have $\max_{a}(\mathcal{B}h)(s,a)<h(s)$
Implication: Maintaining a pessimistic heuristic is inherently fragile because the introduction of a false positive in any state disrupts the pessimistic condition. ∎
See 4.8 We analyze the convergence speed of policy learning by examining the performance gap, which is defined as the difference between the optimal value of the initial state $s_{0}$, $V^{*}(s_{0})$, and the value of the initial state under an arbitrary policy $\pi$, $V^{\pi}(s_{0})$. Specifically, we focus on deriving an upper bound for this performance gap. The key intuition is that a smaller upper bound implies faster convergence to the optimal policy, as fewer iterations of policy updates will be required to reach the optimum.
We begin by stating the theorem made by HuRL authors:
Theorem A.3 (Performance Gap Decomposition (Cheng et al., 2021)).
For any policy $\pi$, heuristic $h:\mathcal{S}\rightarrow\mathbb{R}$, and mixing coefficient $\beta\in[0,1]$,
$V^{*}(s_{0})-V^{\pi}(s_{0})=\mathrm{Regret}(h,\beta,\pi)+\mathrm{Bias}(h,\beta,\pi)$ (30)
where the regret and the bias term are expressed as follows:
$\mathrm{Regret}(h,\beta,\pi)$ $\coloneqq\beta\left(\widetilde{V}^{*}(s_{0})-\widetilde{V}^{\pi}(s_{0})\right)+\frac{1-\beta}{1-\gamma}\left(\widetilde{V}^{*}(d^{\pi})-\widetilde{V}^{\pi}(d^{\pi})\right)$ (31)
$\mathrm{Bias}(h,\beta,\pi)$ $\coloneqq\left(V^{*}(s_{0})-\widetilde{V}^{*}(s_{0})\right)+\frac{\gamma(1-\beta)}{1-\gamma}\mathbb{E}_{s,a\sim d^{\pi}}\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}\left[h(s^{\prime})-\widetilde{V}^{*}(s^{\prime})\right]$ (32)
We will not provide a proof for Theorem A.3 in this paper. Please refer to (Cheng et al., 2021) for details. The theorem demonstrates that the performance gap can be elegantly decomposed into two components: a regret term and a bias term such that:
  1. 1.
    The regret term quantifies the difference between $\widetilde{V^{*}}$ and $\widetilde{V^{\pi}}$, representing the error caused by $\pi$ being suboptimal in the reshaped MDP $\widetilde{\mathcal{M}}$. Since $\pi$ is trained by our selected RL algorithm directly on the reshaped MDP $\widetilde{\mathcal{M}}$, the primary responsibility for minimizing this regret term falls to the RL algorithm itself, not to the design of the auxiliary reward signal. Thus, when evaluating the effects of false positive or false negative rewards, we choose not to focus on bounding the regret term.
  2. 2.
    The bias term captures two key discrepancies: first, between the true optimal value function $V^{*}$ of the original MDP $\mathcal{M}$ and the optimal value function $\widetilde{V}^{*}$ of the reshaped MDP $\widetilde{\mathcal{M}}$; second, between $\widetilde{V}^{*}$ and the heuristic $h(s)$. This term, therefore, reflects the error introduced by addressing the reshaped MDP instead of the original one, alongside how well the heuristic $h$ approximates the optimal value function in the reshaped MDP. Consequently, this bias term directly relates to the quality of the heuristic reward signal $h$. Hence, we focus on analyzing its upper bound to assess the impact of false positive or false negative rewards on this heuristic.
Cheng et al. (2021) have further derived the upper bound for the bias term, which we present here as a lemma. We omit the proof for brevity; for detailed proof, please refer to (Cheng et al., 2021).
Lemma A.4 (Upper Bound of the Bias Term).
$\mathrm{Bias}(h,\beta,\pi)\leq(1-\beta)\gamma\left(\mathbb{E}_{\rho^{\pi^{*}}}\left[\sum_{t=1}^{\infty}(\beta\gamma)^{t-1}(V^{*}(s_{t})-h(s_{t}))\right]+\mathbb{E}_{\rho^{\pi}}\left[\sum_{t=1}^{\infty}\gamma^{t-1}(h(s_{t})-\widetilde{V}^{*}(s_{t}))\right]\right)$ (33)
where $\rho^{\pi}$ denotes the trajectory distribution of $s_{0},a_{0},s_{1},...$ induced by running $\pi$ starting from $s_{0}$.
This upper bound elegantly illustrates a trade-off: if the heuristic $h$ is set too high (overestimation error), it reduces the first term but increases the second term. Conversely, if $h$ is set too low (underestimation error), it increases the first term but reduces the second term. Just by inspection, it is not immediately clear whether overestimation or underestimation results in a better upper bound, as both impact the bias term in opposing ways.
We now prove that underestimation of $h$ (pessimistic $h$) is better than overestimation (i.e., $\exists s_{t}\in\mathcal{S},h(s_{t})>V^{*}(s_{t})$) for minimizing bias.
Proof.
  1. 1.
    Define Terms:
    • Let $B_{1}=\mathbb{E}_{\rho^{\pi^{*}}}\left[\sum_{t=1}^{\infty}(\beta\gamma)^{t-1}(V^{*}(s_{t})-h(s_{t}))\right]$
    • Let $B_{2}=\mathbb{E}_{\rho^{\pi}}\left[\sum_{t=1}^{\infty}\gamma^{t-1}(h(s_{t})-\widetilde{V}^{*}(s_{t}))\right]$
  2. 2.
    Underestimation Case (Pessimistic $h$):
    Lemma A.5.
    If $h$ is pessimistic with respect to $\mathcal{M}$, $\forall\beta\in[0,1],s\in\mathcal{S},\widetilde{V}^{*}(s)\geq h(s)$
    Proof.
    To begin with, we need to use another lemma from Cheng et al. (2021), shown as follows:
    Lemma A.6 (Bellman backup of reshaped MDP (Cheng et al., 2021)).
    For any policy $\pi$, we have
    $\widetilde{V}^{\pi}(s_{0})-h(s_{0})=\frac{1}{1-\beta\gamma}\mathbb{E}_{\widetilde{d}_{s_{0}}^{\pi}}[(\widetilde{\mathcal{B}}h)(s,a)-h(s)]$ (34)
    where $\widetilde{d}_{s_{0}}^{\pi}$ refers to the discounted average state distribution of policy $\pi$ in reshaped MDP $\widetilde{\mathcal{M}}$. It can be expressed as $\widetilde{d}_{s_{0}}^{\pi}=(1-\widetilde{\gamma})\sum_{t=0}^{\infty}\widetilde{\gamma}^{t}d_{t}^{\pi}$, and $d_{t}^{\pi}$ is the state distribution of policy $\pi$ at time $t$ with $d_{0}^{\pi}=\mathbf{1}\{s=s_{0}\}$.
    For brevity, we do not include the proof for Lemma A.6.
    First of all, due to the definition of optimal value $V^{*}$, we have
    $\widetilde{V}^{*}(s_{0})\geq\widetilde{V}^{\pi}(s_{0})$ (35)
    Then, according to Lemma A.6, we can get
    $\widetilde{V}^{*}(s_{0})\geq\widetilde{V}^{\pi}(s_{0})=h(s_{0})+\frac{1}{1-\beta\gamma}\mathbb{E}_{\widetilde{d}_{s_{0}}^{\pi}}[(\widetilde{\mathcal{B}}h)(s,a)-h(s)]$ (36)
    Let $\pi$ denote the greedy policy of $\arg\max_{a}(\mathcal{B}h)(s,a)$ and then trace back to Equation 8, we have (~$(\widetilde{\mathcal{B}}h)(s,a)=(\mathcal{B}h)(s,a)$, that means
    $\widetilde{V}^{*}(s_{0})\geq\widetilde{V}^{\pi}(s_{0})$ $=h(s_{0})+\frac{1}{1-\beta\gamma}\mathbb{E}_{\widetilde{d}_{s_{0}}^{\pi}}[(\widetilde{\mathcal{B}}h)(s,a)-h(s)]$ (37)
    $=h(s_{0})+\frac{1}{1-\beta\gamma}\mathbb{E}_{\widetilde{d}_{s_{0}}^{\pi}}[(\mathcal{B}h)(s,a)-h(s)]$ (38)
    $\geq h(s_{0})\quad\text{[direct result of $h$ being pessimistic]}$] (39)
    Generalization to Any State: The lemma from Cheng et al. (2021) (Lemma A.6) is stated in terms of any starting state $s_{0}$. Therefore, we can replace $s_{0}$ with any state $s\in\mathcal{S}$ in our analysis. Thus we get $\forall s\in\mathcal{S},\widetilde{V}^{*}(s)\geq h(s)$
    Lemma A.5 implies that when $h$ is pessimistic, $B_{2}=\mathbb{E}_{\rho^{\pi}}\left[\sum_{t=1}^{\infty}\gamma^{t-1}(h(s_{t})-\widetilde{V}^{*}(s_{t}))\right]\leq 0$ because $\widetilde{V}^{*}(s)\geq h(s)$ for all $s$. Therefore, the bias error is only bounded by $B_{1}$ when $h$ is pessimistic, i.e., $\mathrm{Bias}(h,\beta,\pi)\leq B_{1}$.
  3. 3.
    Overestimation Case ($\exists s_{t}\in\mathcal{S},h(s_{t})>V^{*}(s_{t})$):
    We show briefly that, unlike the underestimation error, the overestimation error does not have a closed-form upper bound expression. The difficulty, as pointed out by (Cheng et al., 2021), originates from the trajectory-dependence on $\mathbb{E}_{\rho^{\pi}}[\cdot]$ as the $l_{\infty}$ approximation error here can be difficult to control in large state spaces. It is possible that, upon picking up falsely high $h$ states, $\rho^{\pi}$ gets further distorted away from $\rho^{\pi^{*}}$ and accumulates more falsely high $h$ states. In other words, this trajectory dependence makes $B_{2}$ prone to a feedback loop of accumulating overestimation errors and resulting in a much larger upper bound for the bias.
A.3 Additional Notes on Theoretical Setting
A.3.1 Convergence Guarantee of HuRL Compared With Potential-Based Reward Shaping
Previous work on potential-based reward shaping (PBRS) has established that learning a policy on a reshaped MDP can converge to the optimum of the original MDP, as proved by Ng et al. (1999). They proved that the optimal $Q$ value of the reshaped MDP is equivalent to the optimal $Q$ value of the original MDP minus a state-dependent function. Consequently, for an optimal policy defined as $\pi^{*}=\arg\max_{a\in\mathcal{A}}Q^{*}(s,a)$, this additional state-dependent value does not alter the policy. Thus, the optimal policy of the original MDP $\mathcal{M}$ is also the optimal policy for the reshaped $\widetilde{\mathcal{M}}$. However, the theoretical convergence for the reshaped MDP within the framework of Heuristic-Guided Reinforcement Learning (HuRL) has not been proven with the same rigor.
The authors of HuRL made a trick which involves manipulating the coefficient $\beta$, which scales the heuristic-based auxiliary rewards. If $\beta$ increases gradually from 0 to 1 over the training process, the agent effectively transitions from interacting with the modified MDP back to the original MDP. This method ensures that, at the end of training, the agent is exactly solving the original problem. Moreover, according to the Blackwell optimal property (Chen et al., 2018), convergence can occur before $\beta$ reaches 1, thereby allowing the trained policy to maintain optimality in the original MDP under HuRL.
However, in the original HuRL paper, among the five test environments, the $\beta$ hyperparameter was actually fixed for four of them. Surprisingly, HuRL still converged to the optimal policy faster than the original MDP without updating $\beta$. This observation poses an enigma within the HuRL framework, as it is unclear why this convergence to the optimum of the original MDP occurs without dynamically adjusting $\beta$.
Thus, while the practical success of HuRL in these environments is evident, the underlying reasons for such convergence remain to be fully understood. This discrepancy between theoretical expectation and empirical results leaves room for future research to explore why HuRL converges under fixed $\beta$. Nevertheless, proving the convergence properties goes beyond the scope of our paper, which aims to highlight the prevalence of false positive rewards and their impact. An alternative perspective on this issue is to consider the learning process of maximizing auxiliary rewards as akin to performing behavior cloning from experts. In this sense, we do not need to focus on convergence within the current MDP but can view it as a form of pretraining.
A.4 Implementation Details of the Experiments
A.4.1 Environment Details
@@
We describe each testing environment used in our experiments. More details introduction can be found in on the official project homepage of each benchmark (Chevalier-Boisvert et al., 2018; Bellemare et al., 2013; Hafner, 2021).
@@
A.4.2 Instruction-Guide Procedure Details
The VLM-based reward model will have a pointer to the sequence of the instruction sentence, starting at the first sentence. For original models Pixl2R and ELLM-, we follow the setting in there original work where for each instruction sentence (yes, the full instruction essay will be split into multiple sentences and treat each sentence as atomic instruction $l_{k}$), the reward model will have a maximum cap of rewards (2.0) it can assign to the agent in one episode. When the cap is reached, the reward model will move its pointer to the next instruction sentence. For the BiMI reward model, the reward model will move its pointer to the next instruction sentence when the binary signal is triggered. Below is the pseudo-code for the instruction-following RL training procedure for both Pixl2R and ELLM- models.
Algorithm 1 ELLM and Pixl2R pseudo-code for instruction-following RL
1:Initialize policy network $\pi_{\theta}$
2:Initialize value network $V_{\phi}$
3:Setup VLM-based reward model $E(\cdot)$
4:Split instruction essay into sentences $\{l_{1},l_{2},...,l_{K}\}$
5:Initialize instruction pointer $p=1$
6:Initialize cumulative VLM reward $r_{\text{cum}}=0$
7:Initialize cumulative VLM reward threshold $q$
8:Initialize replay buffer $\mathcal{D}$
9:Initialize agent trajectory memory queue $\tau$ with length $W$Foreach episode
[…]
11:Select action $a_{t}\sim\pi_{\theta}(a_{t}|s_{t})$
12:Execute $a_{t}$, observe next state $s_{t+1}$ and extrinsic reward $r^{e}_{t}$
13:Enqueue $(s_{t},a_{t},r_{t},s_{t+1})$ in $\tau$
14:Compute VLM reward:
15:$r^{v}_{t}=\frac{E(\tau)\cdot E(l_{p})}{\|E(\tau)\|\|E(l_{p})\|}$triangleright Apply BiMI here if needed
16:Combine rewards: $r_{t}=r^{e}_{t}+(1-\beta)\gamma r^{v}_{t}$triangleright $\beta$ is a scaling factor
17:Store $(s_{t},a_{t},r_{t},s_{t+1})$ in $\mathcal{D}$
[…]
19:$p\leftarrow\min(p+1,K)$
[…]
21:Sample mini-batch $\{(s_{j},a_{j},r_{j},s_{j+1})\}$ from $\mathcal{D}$
22:Compute TD errors:
23:$\delta_{j}=r_{j}+\gamma V_{\phi}(s_{j+1})-V_{\phi}(s_{j})$
24:Update value network:
25:$\phi\leftarrow\phi+\alpha_{v}\sum_{j}\delta_{j}\nabla_{\phi}V_{\phi}(s_{j})$
26:Update policy network:
[…]
A.4.3 Finetuning VLM-based Reward Models
In contrast to previous work on instructing following RL where they rely on hand-crafted oracle multimodal reward models, we use actual pretrained VLMs to generate reward signals. 2 VLM backbone models are used in our experiments: 1) CLIP (Radford et al., 2021), pretrained by image-text pairs; and (2) X-CLIP (Ma et al., 2022), pretrained by video-text pairs. In particular, Pixl2R uses CLIP because it only uses the single latest frame as input. In contrast, ELLM- takes a slice of trajectory (i.e., multiple frames) as input, and thus uses either X-CLIP or CLIP with additional RNN encoder as the reward model.
@@ A.4.3 Finetuning VLM-based Reward Models
A.4.4 Hyperparameters for Instruction-Following RL Agents
In the experiments, all methods are implemented based on PPO with same model architecture. The Minigrid and Crafter environments use the same training hyperparameters as the Achievement Distillation paper (Moon et al., 2023). For Montezuma’s Revenge, we found that the performance of the agent was sensitive to the gamma and GAE lambda parameters. To improve the performance of agents in Montezuma’s Revenge, we took two additional steps: (1) normalizing the observation inputs when computing the rewards, and (2) not normalizing the advantage during the GAE calculation. The hyperparameters are shown in the following tables.